OpenStudy (anonymous):

Points A(−1, 2), B(−2, −5) and C(7, −2) are on a circle. What is the area of the circle? Express your answer as a decimal to the nearest tenth.

OpenStudy (anonymous):

1. slope of AB 2. mid point of A,B 3. slope of line perpendicular to AB 4. eq. of line (perpendicular bisector) through mid point 5. similarly find the eq. of perpendicular bisector of BC 6. center is point of intersection of perpendicular bisectors (say O) 7.find r=OA 8. area of circle.

OpenStudy (anonymous):

\[slope=\frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\]

OpenStudy (anonymous):

mid point\[=\left( \frac{ x _{1} +x _{2}}{ 2 } ,\frac{ y _{1} +y _{2}}{ 2 }\right)\]

OpenStudy (anonymous):

can you help me i dont know what you are talking about can you give me the answer too @surjithayer

OpenStudy (anonymous):

first find the slope of AB by the above formula and show me.

OpenStudy (anonymous):

\[m=\frac{ -5-2 }{ -2-(-1) }=\frac{ -7 }{ -2+1 }=\frac{ -7 }{ -1 }=7\] slope of line perpendicular to AB=-1/m=-1/7

OpenStudy (anonymous):

mid point of A,B

OpenStudy (anonymous):

will you show me ?